

********************************************************************
UARS WEEKLY STATUS REPORT - 03 Sep 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 56% of
each day's data amounting to nearly 525 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004099 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004089. Currently, MEPS data between
2004090 and 2004099 has been processed waiting for archival at the GSFC DAAC.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothling received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 August 2004 (Orbit 70834, GMT Day 234) through
27 August 2004 (Orbit 70939, GMT Day 240).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 234/1233 70842 DAY ONLY TO FULL TIME
SUSIM 234/1453 70843 POWER ON
HALOE 236/2246 70878 POWER ON
PEM 237/1338 70888 ZEPS/NEPS DAY-ONLY
PEM 237/1543 70889 AXIS 1 & 2 ON
The beta angle ranged between the angles of 11.2 degrees to 35.1
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 63 degrees on 06 September 2004 (DOY 250).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.96 degrees C. Additional
battery performance data for each battery on 13 August and 20 August
2004 are provided in Appendix D.
The clock error ranged from 08.1 to -07.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
SOLSTICE = 01
SUSIM = 01
HALOE = 02
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On 10 August 2004, during the 223/0121 Z event, VT-6 was commanded due
to the differential voltage being greater than 180 mV for three
consecutive orbits. The differential voltage returned to nominal
values after two orbits subsequent to the change. The spacecraft
remains in VT-6 and will be monitored closely to determine when to go
back to VT-5.
The new CMS code to allow for SMA events continued to undergo testing
this week. Problems receiving products from FDF, transferring the
schedule from UPS to CMS, and system drive issues hindered progress.
While the FDF product and system drive issues have been resolved, the
transfer issue is still under investigation.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 25 September 2004 (DOY 269).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 August 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70842 234/1233 04-240 SOLSTICE DAY ONLY TO FULLTIME
70843 234/1453 04-241 SUSIM POWER ON (PART 1)
70849 235/1050 04-241 SUSIM POWER ON (COMPLETE)
70875 236/1729 04-242 HALOE POWER ON (WARMUP)
70878 236/2246 04-243 HALOE POWER ON
70888 237/1338 04-244 PEM ZEPS/NEPS DAY ONLY
70889 237/1543 04-245 PEM AXIS 1 & 2 ON
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 August 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 August 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 August 2004
20 August 2004 - GMT Day 233 Beta = 07.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-044.8 +5.15/+4.19 +27.2/-12.0 26.4 1.04 18.8
3 +00.0/-112.0 +3.87/+2.92 +25.6/-12.0 26.4 1.05 18.2
27 August 2004 - GMT Day 240 Beta = 35.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.73/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-050.4 +5.79/+4.83 +26.8/-14.4 26.4 1.02 19.7
3 +00.0/-100.8 +4.51/+3.87 +25.6/-14.0 26.4 1.02 19.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 August 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-057 FOT 70751 228/1124 36677 Drop out on the Return
04-058 FOT 70774 230/0059 36694 7:01 Data loss:
weather@WSGT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-050 FOT See List August 04 30832 Generic Late Acquisition